OpenStudy (anonymous):

Jordan invested $1000 in a savings account. The interest rate is 6% per year. Find the simple interest earned in 4 years. Then find the total of principal plus interest

OpenStudy (anonymous):

Do you know the formula to calculate simple interest?

OpenStudy (anonymous):

I = P x r x T

OpenStudy (anonymous):

\[I=PRT\]\[P=1,000\]\[R=0.06\]\[T=4\]\[=1,000*0.06*4\]Can you solve it?

OpenStudy (anonymous):

\[1,000*0.06*4 =\]\[60*4=240\]240 is the answer ;)